OpenStudy (anonymous):

Algebra question? 8-wx=vx+2 what does x=

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

8-2 = vx + wx 6=x(v+w) x = 6/(v+w)

OpenStudy (anonymous):

thanks him1618 solved